OpenStudy (anonymous):

Simplify the expression below. 169

OpenStudy (anonymous):

\[\sqrt{169}\]

OpenStudy (mathstudent55):

1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 ... 10^2 = 100 11^2 = 121 12^2 = 144 13^2 = 169 14^2 = 196 etc.

OpenStudy (anonymous):

so is that the answer

OpenStudy (mathstudent55):

The answer is 13. Square root is the opposite of squaring. If 13 squared equals 169, then the square root of 169 equals 13